Experience has shown that the time taken to drive from city A to city B is normally distributed, with an average time (µ) of 65 minutes, and a standard deviation (σ) of 11 minutes. A sample of 88 driving trips between the two cities is randomly selected.  Determine the standard error for this sample.


1
Expert's answer
2021-10-21T12:00:08-0400
"SE=\dfrac{\sigma}{\sqrt{n}}=\dfrac{11}{\sqrt{88}}\approx1.1726"